Shift summary for day shift, 01/04/2000


Subject: Shift summary for day shift, 01/04/2000
From: Vivian O'Dell (odell@fnal.gov)
Date: Wed Jan 12 2000 - 16:38:36 CST


        Shift Summary for Day Shift, 01/04/2000

On Shift: E. Cheu, R. Coleman
Hours of beam available at all intensities: 3.0
Hours of beam available at nominal intensity: 3.0
Hours of E799 data (special runs not included): 3.0
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:

Booster and MI power outtage, followed by MI ground fault investigation
========================================================================
Runs ended this shift:
Run 15390 -- BEAM_799 11.0 M events

Tapes brought to FCC this shift: none
  (including tapes for online split)

========================================================================
Accesses during shift: Access for Bob H to gather more phototubes for
upcoming KAMI test

========================================================================
Overview of shift:

E799 data taking first 3 hours
Beam down due to Booster power outtage (brief), MI power outtage(longer)
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15390 07:36 4x-1 Kumquat module was not reading out,
                        Jason L. arrives and finds loose barrel connection
                        in L1 signal going to Kumquats, 4x-1 fixed.
 
                        
               10:30 Booster loses power.

               10:50 Main Injector has a power outtage also.
                        After power is restored, a Main Injection
                        ground fault is present, investigating....

               12:15 Leo notices 1-Y DC OR ineff. 0.999 has dropped
                        to 0.993, Ashkan is looking into it.

               14:20 Magnets set for muon run when beam returns,
                        Mike A. has installed new BA voltages for
                        muon run, need to return to nominal after muon run.

               15:00 MCR continues Main Injector ground fault.
                        Next estimate 16:30.

========================================================================
Items to Follow up on Next Shift

                        Mike A. has installed new BA voltages for
                        muon run, need to return to nominal after muon run.



This archive was generated by hypermail 2b27 : Wed Jan 12 2000 - 16:38:39 CST